Scott’s workout

• Sunday

– Quads ∗ Squats ( 5 reps min/set 6 sets )

∗ Leg Press ( 5 reps min/set 6 sets )

∗ Alternate Extensions,Hack Squats, etc

– Chest ∗ Bench 5 reps min, 5-6 sets)

∗ Alternate Incline Bench, Dumbells 5 reps min, 5-6 sets)

∗ Alternate (Flyes, Incline Flyes, Machine Flyes, Cables) 4 sets – Treadmill 0.5 hr

• Monday

– Shoulders ∗ Alternate Dumbell presses/smith rack (5 rep min/ 6 sets )

∗ Alternate Laterals/(Front movement) (8 rep min/ 3 sets )

∗ Alternate Upright rows/Front presses (8 rep min/ 3 sets)

∗ Alternate Shrugs/Rear machine presses (8 rep min/ 3-4 sets)

– Biceps ∗ Standing Curls ( 5 rep min/5 sets)

∗ Alternate Preacher/rotating curls ( 5 rep min/4 sets)

∗ Alternate Conc. curls/cable curls/etc. ( 4 sets)

– Treadmill 0.5 hr

• Wenesday

– Deadlifts/hamstrings ∗ Dead Lifts (5 rep min/ 6 sets)

∗ Standing curls

∗ lying curls

– Back ∗ Wide grip chinups 3 sets/ 10 reps

∗ Pull downs (alt. wide/med. narrow) 4 sets/ 6-8 reps

∗ Alt TBar/ lying T bar 4 sets, vary grip (6-8 reps )

∗ Rows ( 3-4 sets, 8-10 reps)

– Abs

• Thursday

– Triceps ∗ Presses ( 4 sets, 10 reps)

∗ Alternate Skull Crunchers,Close Grip Bench (4 sets, 6-10 reps)

∗ Alternate Rev. Skull Rev Presses, Rope ( 4 sets 8-10 reps)

– Calves ∗ Standing Calf raises

∗ Seated Calf raises

– Abs

– Treadmill 40 min